### v1

# Enhancing Chemistry on Quantum Computers with Fermionic Linear Optical Simulation

Abstract
We present and open source a simulator for circuits composed of passive fermionic linear optical elements and controlled-phase gates. Given such a circuit, our simulator can compute Born-rule probabilities for samples drawn from it. Our simulator supports both exact and approximate probability calculation, allowing users to trade accuracy for efficiency as needed. For approximate Born-rule probability calculation, our simulator’s runtime is exponential only in the magnitudes of the angles of the circuit’s controlled-phase gates. This makes our simulator useful for simulating certain systems that are beyond the reach of conventional state vector methods. We demonstrate our simulator’s utility by simulating the local unitary cluster Jastrow (LUCJ) ansatz and integrating it with sample-based quantum diagonalization (SQD) to improve the accuracy of molecular ground-state energy estimates. Applied to a 52-qubit N2 system, we observe accuracy improvements of up to 46% over the baseline SQD implementation with negligible computational overhead. As an efficient and flexible tool for simulating fermionic circuits, our simulator enables new opportunities for enhancing near-term quantum algorithms in chemistry and related domains.
I Introduction
Among the most promising near-term applications of quantum computing is quantum computational chemistry [cao2019quantum, mcardle2020quantum], which has attracted considerable attention due to potentially transformative applications in drug development [santagati2024drug] and materials science [alexeev2024quantum], among other fields. Despite the excitement around this application, hardware demonstrations have been limited by the noise present in current quantum processors. Quantum error correction [gottesman1997stabilizer, fowler2012surface, shor1995scheme, steane1996error, calderbank1996good, kitaev2003fault] offers a long-term solution to faulty hardware, but a fully fault-tolerant quantum computer has yet to be built. In the near-term, quantum error mitigation techniques [temme2017error, li2017efficient, huggins2021virtual, liao2024machine, cincio2021machine] can improve results obtained from noisy quantum processors, in which many of these techniques involve classical simulation of quantum circuits [czarnik2021error, filippov2023scalabletensornetworkerrormitigation, eddins2024lightcone]. Because classical simulation of quantum circuits is intractable in general, these simulations must introduce approximations, or are exact only when applied to specific narrow classes of circuits. Simulation methods tailored for circuits relevant to quantum chemistry may enable new techniques for enhancing the performance of chemistry workloads on quantum computers.
In this work, we introduce ExtraFerm [^1] https://github.com/zhassman/ExtraFerm, an open-source quantum circuit simulator that can be used to compute the Born-rule probabilities of samples (bitstrings) from circuits composed of passive fermionic linear optical elements and controlled-phase gates. These circuits are highly relevant to the field of quantum computational chemistry, in which passive fermionic linear optical elements are also known as particle number-conserving matchgates. ExtraFerm can perform both exact and approximate probability calculations for measurements in the output distributions of these circuits. For approximate calculation, the only exponential component of ExtraFerm’s runtime is a linear dependence in a quantity called the extent of the circuit. The extent is larger for controlled phase gates with larger phases and is multiplicative as more controlled phase gates are added (see Refs. [reardon2024improved, Dias2024classicalsimulation, reardonsmith2024fermioniclinearopticalextent]). For exact probability calculation, the runtime is exponential only in the number of controlled-phase gates.
Notably, this means that for both approximate and exact calculation, ExtraFerm’s performance is polynomial in the number of qubits and matchgates. Unlike conventional state vector simulation, which tracks the amplitudes of all bitstrings, ExtraFerm computes Born-rule probabilities for a pre-specified subset of the output distribution. Indeed, for small systems, one can use ExtraFerm to simulate the entire support of a circuit, but we do not expect this to be the most compelling use case. Instead, we propose ExtraFerm as a unique tool that can be embedded in quantum-classical workflows to recover signal from noisy samples from large, application-scale quantum circuits.
We apply ExtraFerm to the simulation of the local unitary cluster Jastrow (LUCJ) ansatz [motta2023lucj], which has recently been adopted for diverse applications in quantum simulation of chemical systems [blunt2025quantum, shajan2025toward, kaliakin2025implicit_solvent, liepuoniute2025quantumcentric, robledo2025sqd]. When mapped to a quantum circuit using the Jordan-Wigner transformation [ortiz2001quantum], the LUCJ ansatz decomposes into particle number-conserving matchgates and controlled-phase gates, making it amenable to simulation by ExtraFerm in certain regimes. To highlight the regimes in which ExtraFerm may be useful, we use LUCJ circuits to study the error and latency performance of ExtraFerm. Note that LUCJ circuits are flexible in the sense that repeated layers of them can be used to implement arbitrary particle number-conserving extended matchgate circuits, providing us with a generic but practically useful benchmark.
Finally, as an end-to-end practical use case, we show how to use ExtraFerm to boost the performance of sample-based quantum diagonalization (SQD) [robledo2025sqd]. SQD is an extension of quantum-selected subspace interaction (QSCI) [kanno2023quantum], an algorithm that samples configurations from a quantum computer and uses them to select a subspace in which to diagonalize a molecular Hamiltonian. SQD adds an error mitigation procedure called configuration recovery that attempts to correct sampled bitstrings that were affected by noise and do not satisfy symmetries of the system. Configuration recovery greatly improves performance on noisy quantum processors, leading to the adoption of SQD for some of the largest demonstrations of quantum chemistry on quantum processors to date [liepuoniute2025quantumcentric, kaliakin2025implicit_solvent, shajan2025toward].
We improve SQD by using ExtraFerm to select high-probability bitstrings during the early iterations of configuration recovery to provide better orbital occupancy information. We call this approach the “warm-start” variant of SQD and find that, compared to the original implementation of SQD, we obtain higher-accuracy energy estimates. For our main result, we consider a 52-qubit system sampled on an IBM Heron quantum processor and observe accuracy improvements of up to 46.09% with respect to the heat-bath configuration interaction (HCI) energy taken as a reference. Across repeated trials, we find that the warm-start variant enjoys a variance reduction of up to 98.34%. Furthermore, the overhead of the warm-start variant is minimal—at worst accounting for an extra 2.03% of runtime. For completeness, we also report results for a 28-qubit system simulated classically in the presence of global depolarizing noise and observe similar improvements. The main contributions of this work are the following:
  1. 1.
    We release ExtraFerm, an extended matchgate simulator for particle number-conserving circuits allowing for high-performance calculation of Born-rule probabilities with tunable accuracy.
  2. 2.
    We evaluate latency and error performance of ExtraFerm on local unitary cluster Jastrow (LUCJ) circuits to identify regimes for which ExtraFerm excels.
  3. 3.
    We demonstrate how to integrate our ExtraFerm with SQD, a hybrid quantum-classical energy estimation algorithm to improve the accuracy and variance of molecular ground-state energy calculations.
Figure 1: A high-level depiction of how ExtraFerm implements the Raw Estimate and Exact algorithms. The number of trajectories, $t$, can be given arbitrarily for Raw Estimate, while Exact sums over all possible trajectories. In practice, we parallelize this computation across bitstrings and then parallelize again across trajectories. Separate trajectories are generated for each bitstring to preserve the independence of Born-rule probability estimates.
II Background
II-A Classical Simulation of Quantum Circuits
It is important to distinguish the very different computational tasks which are termed “simulation”. ExtraFerm computes additive-precision estimates of individual Born-rule probabilities and can be thought of as an approximate strong simulator in the sense of Ref. [jozsa2013classical]. Conversely, a weak simulator would sample from the output distribution of a quantum circuit. While it might seem intuitive that a strong simulator is stronger than a weak simulator, the precise comparison between the different simulation tasks is subtle and depends strongly on the error allowed [pashayan2020estimation]. However, our simulator can be used to approximately sample from the Born-rule distribution when combined with the Qubit-by-qubit algorithm [bravyi2022simulate] or the Gate-by-gate algorithm.
An advantage of our methods is the ability to target specific Born-rule probabilities, while a state vector simulator computes all $2^{n}$ Born-rule probabilities for an $n$-qubit system. In applications such as the one we consider in Section VI, we are only interested in a small subset of the output distribution. As we will later discuss, the fact that we can use ExtraFerm to simulate arbitrary subsets of the output distribution allows us to improve SQD with very little computational overhead.
II-B The Theory of Matchgates
Matchgates are a class of classically simulable quantum gates formalized by Valiant [valiant2002quantum]. Shortly afterwards, Terhal and DiVincenzo [terhal2002classical] discovered that matchgates are an equivalent expression of non-interacting fermions, giving these circuits a natural physical interpretation. The term ‘extended matchgate’ refers to circuits that are primarily composed of matchgates but include a limited quantity of non-matchgates—in our case, controlled-phase gates. Together, matchgates and controlled-phase gates allow for universal quantum computation [PhysRevA.73.042313, brod2011extending]. However, current matchgate-based chemistry ansatze are particle number-conserving, meaning that the Hamming weight of all measurement outcomes are the same as the initial Hamming weight of the system, as specified by Pauli-X gates. These are the circuits that ExtraFerm is designed for.
More formally, a matchgate is a 2-qubit gate that has a $4\times 4$ unitary matrix of the form
$G(A,B)=\begin{pmatrix}a_{11}&0&0&a_{12}\\ 0&b_{11}&b_{12}&0\\ 0&b_{21}&b_{22}&0\\ a_{21}&0&0&a_{22}\end{pmatrix}$ (1)
such that $\det(A)=\det(B)$ [jozsa2008matchgates]. Via the Jordan-Wigner transformation [ortiz2001quantum] we associate an $n$ qubit state with an $n$ mode fermionic state. A qubit state with qubit $i$ in state $\left|1\right\rangle$ maps to a fermionic state with a fermion in mode $i$. Let $M$ be a circuit of nearest-neighbor, particle number-conserving matchgates. Such a circuit preserves the Hamming weight of the computational basis state that it is applied to and is also sometimes called ‘passive’ in the literature. The action of $M$ can be described by a unitary matrix $V$ which satisfies
$Ma_{i}^{\dagger}M^{\dagger}=\sum_{j}V_{ij}a_{j}^{\dagger}$ (2)
where $a_{i}^{\dagger}$ denotes the $i^{\text{th}}$ creation operator defined via the Jordan-Wigner transformation. Note that while $M$ is a $2^{n}\times 2^{n}$ unitary matrix, $V$ is only $n\times n$. We refer to these ‘$V$’ matrices as mode transformation matrices. Let $\left\langle b\right|$ and $\left|a\right\rangle$ be two computational basis states with the same Hamming weights. Let $R$ be the indices of the bits of $\left\langle b\right|$ that are equal to 1 and let $C$ be the indices of the bits of $\left|a\right\rangle$ that are equal to 1.
[…]
where $\widetilde{V}$ is the submatrix of $V_{\text{circuit}}$ obtained by selecting rows $r\in R$ and $c\in C$ [terhal2002classical].
II-C Simulation of Extended Matchgate Circuits
ExtraFerm builds upon the recently proposed Algorithm 1 from Reardon-Smith et al. [reardon2024improved] and optimizes it for particle number-conserving extended matchgate circuits. We provide the details of this in section III-A. Simultaneously with [reardon2024improved], Diaz and Koenig [Dias2024classicalsimulation] developed alternative methods for extended matchgate simulation with essentially identical runtime scaling. Very recently, another approach has been developed with worse asymptotic scaling but significant improvements to the polynomial component of the runtime [wille2025classicalsimulationparitypreservingquantum]. These polynomial improvements may prove more relevant than the exponential component of the runtime in some parameter regimes.
Finally, approaches involving decompositions of operators rather than state vectors have also been explored, e.g. in Refs [mocherla2023extending, Hakkaku_2022], but these suffer from a substantially worse asymptotic scaling than state vector methods. For all of these methods, the runtime is dominated by an exponential sampling cost, but operator decomposition methods generally require a quadratic number of samples, i.e. the exponentially large component of the runtime is squared. Crucially, while there are many proposed mathematical frameworks for simulating extended matchgate circuits, there are few tools that exist for doing this in practice. ExtraFerm aims to contribute to filling this gap for particle number-conserving circuits.
III Simulator
III-A Algorithms
ExtraFerm is capable of performing three different algorithms: Raw Estimate, Estimate, and Exact. Raw Estimate and Estimate, whose names follow the formalization in Ref. [pashayan2022fast], compute estimates of Born-rule probabilities, while Exact computes probabilities exactly. The Raw Estimate algorithm, which is depicted in Fig. 1, is the flagship capability of ExtraFerm. Given an additive error $\epsilon$, failure probability $\delta$, and probability upper bound $p_{\text{max}}$, Raw Estimate estimates the Born-rule probability $\hat{p}$ of a bitstring such that
$\mathbb{P}\left[\left|\hat{p}-p_{\text{true}}\right|>\epsilon\right]=1-\delta.$ (4)
Raw Estimate does this by summing up contributions from trajectories, which are classically simulable approximations of the original circuit. Intuitively, each time we encounter a controlled-phase gate, we probabilistically branch between two possible paths. The number of trajectories that are needed to achieve a Born-rule probability estimate within $\epsilon$ for failure probability $\delta$ is exponential only in the extent $\xi^{\ast}$ of the circuit, given by the formula
$\xi^{\ast}=\prod_{j=1}^{k}\left(\cos\left(\frac{|\theta_{j}|}{4}\right)+\sin\left(\frac{|\theta_{j}|}{4}\right)\right)^{2}.$ (5)
where $\theta_{j}$ is the angle of the $j^{th}$ controlled-phase gate in the circuit. Using the parameters $\epsilon$, $\delta$, $p_{\text{max}}$, and $\xi^{\ast}$, a lower bound for the number of required trajectories is given by
$t=\left\lceil 2\frac{\left(\sqrt{\xi^{*}}+\sqrt{p_{\max}}\right)^{2}}{\left(\sqrt{p_{\max}+\epsilon}-\sqrt{p_{\max}}\right)^{2}}\log\left(\frac{2e^{2}}{\delta}\right)\right\rceil.$ (6)
In practice, one may not always have particular error bounds in mind. To allow for heuristic use of ExtraFerm, users can directly provide an arbitrary trajectory count $t$ instead of $(\epsilon,\delta,p_{\text{max}})$. Furthermore, it is often not clear a priori what the associated $p_{\text{max}}$ is for a measurement outcome. Indeed, one can always provide $p_{\text{max}}=1$, but this can lead to an unnecessarily large trajectory count, especially for small $\epsilon$ and $\delta$. In this case, one can use ExtraFerm’ Estimate algorithm. This algorithm only requires $(\epsilon,\delta)$ as input parameters, and estimates Born-rule probabilities by starting with loose error and probability upper bounds followed by iterative calls to Raw Estimate. Finally, ExtraFerm can also compute exact probabilities with the Exact algorithm, which simply sums the contributions over all possible trajectories rather than just over the lower bound given by equation 6.
III-B Generating Trajectories
Figure 2: A visualization of how a trajectory $x$ and corresponding mode transformation matrix $V(x)$ is generated from a circuit. Each matchgate $m_{i}$ is unchanged while the controlled-phase gates are probabilistically assigned to either $d_{0}(\theta_{j})$ or $d_{0}(\theta_{j})$ based on $|\theta_{j}|$. The first controlled-phase gate is assigned to $d_{0}(\theta_{0})$ with probability 11 while the second controlled-phase gate is assigned to $d_{0}(\theta_{1})$ with probability $1/2$. Note that controlled-phase gates may operate on any two qubits and are not restricted to adjacent qubits.
Consider a circuit composed of particle number-conserving matchgates and $k$ controlled-phase gates. Our ability to generate trajectories relies upon the following observation: notice that the $j^{th}$ controlled-phase gate $c(\theta_{j})$ can be decomposed into a weighted sum of two unitary matrices, $d_{0}(\theta_{j})$ and $d_{1}(\theta_{j})$
$c(\theta_{j})$ $=e^{i\frac{\theta_{j}}{4}}\Bigg[\cos\!\left(\tfrac{\theta_{j}}{4}\right)\underset{d_{0}(\theta_{j})}{\underbrace{\begin{pmatrix}e^{-i\frac{\theta_{j}}{2}}&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&e^{i\frac{\theta_{j}}{2}}\end{pmatrix}}}$
$\quad+i\sin\!\left(\tfrac{\theta_{j}}{4}\right)\underset{d_{1}(\theta_{j})}{\underbrace{\begin{pmatrix}e^{-i\frac{\theta_{j}}{2}}&0&0&0\\ 0&-1&0&0\\ 0&0&-1&0\\ 0&0&0&e^{i\frac{\theta_{j}}{2}}\end{pmatrix}}}\Bigg$] (7)
One can check that $d_{0}(\theta_{j})$ and $d_{1}(\theta_{j})$ are both particle number-conserving matchgates. For a circuit containing $k$ controlled-phase gates, there are $2^{k}$ ways to replace the controlled-phase gates with either $d_{0}(\theta_{j})$ or $d_{1}(\theta_{j})$. We refer to these sequences as trajectories $x\in\{0,1\}^{k}$, for which the $j$th component $x_{j}$ indicates the choice for gate $d_{x_{j}}(\theta_{j})$. We obtain the exact Born-rule probability $p_{\text{true}}$ with
$p_{\text{true}}=\left|\sum_{x\in\{0,1\}^{k}}w(x)\left\langle b\right|V(x)\left|a\right\rangle\right|^{2}$ (8)
where
$w(x)=\prod_{j=1}^{k}\cos\left(\frac{\theta_{j}}{4}\right)^{x_{j}-1}\cdot\left(i\sin\left(\frac{\theta_{j}}{4}\right)\right)^{x_{j}}$ (9)
This calculations is the basis of the Exact algorithm (see Appendix section -A for derivation).
Allowing for a small additive error $\epsilon$ and failure probability $\delta$, Monte Carlo methods can be used to obtain an estimate for this Born-rule probability. Rather than summing up the amplitudes contributed by all $2^{k}$ trajectories, we can generate sufficiently many trajectories from the product probability distribution $P=\prod_{j=1}^{k}p(x_{j})$:
$p(x_{j}=0)$ $\equiv\frac{\cos\left(\frac{|\theta_{j}|}{4}\right)}{\sin\left(\frac{|\theta_{j}|}{4}\right)+\cos\left(\frac{|\theta_{j}|}{4}\right)}$ (10)
$p(x_{j}=1)$ $\equiv\frac{\sin\left(\frac{|\theta_{j}|}{4}\right)}{\sin\left(\frac{|\theta_{j}|}{4}\right)+\cos\left(\frac{|\theta_{j}|}{4}\right)}$ (11)
To generate a trajectory, for every $\theta_{j}$, one can draw $u_{j}\sim\text{Unif}[0,1]$ and set
$x_{j}=\begin{cases}1,&u_{j}<\dfrac{\sin\left(|\theta_{j}|/4\right)}{\sin\left(|\theta_{j}|/4\right)+\cos\left(|\theta_{j}|/4\right)}\\ 0,&\text{otherwise}\end{cases}$ (12)
Observe that when $\theta_{j}=0$, we select $d_{0}(\theta_{j})$ with probability 1. Conversely, the worst case occurs when $\theta_{j}=\pi$ and we have an equal chance of selecting either $d_{0}(\theta_{j})$ or $d_{1}(\theta_{j})$. A visualization of this is provided by Fig. 2.
Algorithm 1 Raw Estimate
1:Input:
  • computational basis states $\left\langle b\right|$ and $\left|a\right\rangle$
  • quantum circuit $Q$ (consisting of passive matchgates and controlled-phase gates)
  • trajectory count $t$ (Eq. 6)
  • circuit extent $\xi^{\ast}$ (Eq. 5)
2:Output: $\hat{p}\approx\left|\left\langle b\right|Q\left|a\right\rangle\right|^{2}$
[…]
4:for $t$ iterations do
5:  generate trajectory $x\sim P$
6:  construct $V(x)$
7:  $z\leftarrow\left\langle b\right|V(x)\left|a\right\rangle$
8:  construct negative controlled-phase angle mask $m$
9:  calculate sign $s\leftarrow(-1)^{|m\land x|}$
10:  accumulate $\alpha\leftarrow\alpha+i^{|x|}\cdot z\cdot s$
11:end for
[…]
13:return $\hat{p}$
Algorithm 2 Estimate
1:Input:
  • computational basis states $\left\langle b\right|$ and $\left|a\right\rangle$
  • quantum circuit $Q$ (consisting of passive matchgates and controlled-phase gates)
  • circuit extent $\xi^{\ast}$ (Eq. 5)
  • additive error $\epsilon$
  • failure probability $\delta$
2:Output: $\hat{p}\approx\left|\left\langle b\right|Q\left|a\right\rangle\right|^{2}$
[…]
6:while $\epsilon^{\ast}>\epsilon$ do
7:  $\hat{p}\leftarrow\textsc{Raw Estimate}(\left\langle b\right|,\left|a\right\rangle,Q,\xi^{\ast},t)$
8:  $p^{*}\leftarrow\mathrm{clamp}\left(\min(p^{*},\,\hat{p}+e^{*}),\,0,\,1\right)$
9:  increment $k$
10:  $\delta_{k}\leftarrow\frac{6\delta}{\pi^{2}k^{2}}$
11:  double $t$
12:end while
13:return $\hat{p}$
IV Implementation Details
IV-A Overview
We have open sourced ExtraFerm on Github, and it can be installed with pip install extraferm. ExtraFerm consists of a Python interface and Rust backend. The Python interface integrates directly with Qiskit [qiskit], while the Rust backend allows for multithreading of compiled binaries across multiple CPU cores, when available. Interaction between Python and Rust is facilitated by the PyO3 library, which allows Rust functions to be directly called directly from Python.
The machinery behind ExtraFerm’s Raw Estimate and Exact algorithms, depicted in Fig. 1, is doubly parallel. We use Rust’s Rayon library to parallelize the per-bitstring probability computations as well as each bitstring’s per-trajectory computations. A variety of other small optimizations led to further performance increases. For example, we represent states in a bit-packed format: Python uses native integers, while Rust leverages unsigned 128-bit integers (u128), enabling efficient bitwise operations. Additionally, we precompute and store values of $\sin(\theta/4)$ and $\cos(\theta/4)$ for all $\theta$, which reduces the number of computations inside hot loops.
IV-B LUCJ-specific Optimizations
Due to the growing popularity of the LUCJ ansatz, we equip ExtraFerm with specific optimizations for these circuits. LUCJ circuits, which we will denote with $\left|\Psi\right\rangle$ have the form
$\left|\Psi\right\rangle=e^{-\hat{K}_{2}}e^{\hat{K}_{1}}e^{i\hat{J}_{1}}e^{-\hat{K}_{1}}\left|\mathbf{x_{\text{RHF}}}\right\rangle,$ (13)
for which $e^{-\hat{K}_{2}},e^{\hat{K}_{1}},e^{-\hat{K}_{1}}$ are orbital rotations, $e^{i\hat{J}_{1}}$ is a cluster operator, and $\left|\mathbf{x_{\text{RHF}}}\right\rangle$ is the restricted Hartree-Fock (RHF) state as described by the Jordan-Wigner transformation. The orbital rotation operators $e^{-\hat{K}_{2}},e^{\hat{K}_{1}},e^{-\hat{K}_{1}}$ of $\left|\Psi\right\rangle$ can be fully decomposed into particle number-conserving matchgates. On the other hand, the operator $e^{i\hat{J}_{1}}$ can be fully decomposed into controlled-phase gates.
Let $V_{2}$ be the mode transformation matrix describing the action of $e^{-\hat{K}_{2}},e^{\hat{K}_{1}}$ and $V_{1}$ be the mode transformation matrix describing the action of $e^{-\hat{K}_{1}}$. Now let $D_{x_{j}}(\theta_{j})$ be the mode transformation matrix representing the action of $d_{x_{j}}(\theta_{j})$. Our circuit can be written with respect to a trajectory $x$ as
$V(x)=V_{2}\left(\prod_{j=1}^{k}D_{x_{j}}(\theta_{j})\right)V_{1}$ (14)
It is clear that $D_{0}^{-1}\cdot D_{0}=I$, and one can check that
$D_{0}^{-1}(\theta_{j})\cdot D_{1}(\theta_{j})=\text{Diag}\bigl(1,\ldots,\underbrace{-1}_{q1},\ldots,\underbrace{-1}_{q2},\ldots,1\bigr)$ (15)
where $q_{1},q_{2}$ are the two qubits that $D_{x_{j}}(\theta_{j})$ acts on. Notice that $W\equiv\prod_{j=1}^{k}D_{0}^{-1}(\theta_{j})\cdot D_{x_{j}}(\theta_{j})$ is a diagonal matrix with only $\{+1,-1\}$ on its diagonal. Notably, $W$ has no dependence on $\{\theta_{j}\}$. Define $V_{3}=V_{2}\cdot\prod_{j=1}^{k}D_{0}(\theta_{j})$ and reformulate $V(x)$ as
$V(x)$ $=V_{2}\left(\prod_{j=1}^{k}D_{0}(\theta_{j})\cdot D_{0}^{-1}(\theta_{j})\cdot D_{x_{j}}(\theta_{j})\right)V_{1}$ (16)
$=V_{3}\left(\prod_{j=1}^{k}D_{0}^{-1}(\theta_{j})\cdot D_{x_{j}}(\theta_{j})\right)V_{1}$ (17)
$=V_{3}\left(I-2\cdot\sum_{i\in N}E_{ii}\right)V_{1}$ (18)
$=V_{3}V_{1}-2\cdot\sum_{i\in N}V_{3}E_{ii}V_{1}$ (19)
where $N=\{i|W_{ii}=-1\}$ and $E_{ii}$ is a standard basis matrix. In this sense, any possible $V(x)$ can be written as a base $V_{3}V_{1}$ with a small number of corrections of the form $V_{3}E_{ii}V_{1}$ applied to it. ExtraFerm computes all of these matrices once at runtime. Additionally, ExtraFerm stores the determinants corresponding to each unique pattern of corrections to the base matrix in a hash table to avoid redundant determinant evaluations across trajectories.
These optimizations greatly reduce the two primary costs of Born-rule probability computation, namely, the construction of mode transformation matrices $V(x)$ and the evaluation of their determinants. We find that these optimizations lead to speedups of up to two orders of magnitude for LUCJ circuits over our standard implementation of Raw Estimate. When ExtraFerm detects an LUCJ circuit, this optimized path is automatically taken.
V Benchmarking and Error Analysis
V-A Comparisons with Existing Tools
Figure 3: Latency and memory comparisons of different tools when used to calculate the exact probability of an outcome measurement from an arbitrary particle number-conserving extended matchgate circuit. Each circuit contains 16 controlled-phase gates, and the quantity of matchgates grows with circuit size. Results are collected on an Intel Xeon w7-2495X processor. By default, all tools are given access to 24 cores, but only ExtraFerm benefits significantly from using more than 1. We also include results for ExtraFerm restricted to 1 core. Note that the y-axis for each graph is on the logarithmic scale (higher values are orders of magnitude worse).
We emphasize that while LUCJ circuits are an important use case for ExtraFerm, our simulator is a competitive tool for simulating particle number-conserving extended matchgate circuits in general. To demonstrate this, we compared ExtraFerm’s latency and memory consumption against ffsim [ffsim], a state vector simulator that takes advantage of particle number symmetry, and quimb [gray2018quimb], a tensor network simulator. While ffsim computes probabilities for the entire output distribution, quimb is similar to ExtraFerm in the sense that it can compute the probability for a single outcome measurement. We benchmark these three tools on randomly-initialized particle number-conserving extended matchgate circuits ranging from 8 to 36 qubits. These circuits contain 16 controlled-phase gates and a quantity of matchgates proportional to the circuit’s size. Note that these more general circuits are not amenable to the optimizations for LUCJ circuits described in Section IV-B and thus they are not applied.
Fig. 3 demonstrates that for ExtraFerm, latency increases linearly with respect to the number of qubits in the circuit, while for ffsim and quimb, latency increases exponentially. ExtraFerm’s embarrassingly parallel implementation allows it to take full advantage of all 24 CPU cores. On the other hand, ffsim and quimb do not significantly benefit from using more than one core. In terms of memory consumption, system size hardly has an effect on ExtraFerm, while quimb and ffsim suffer an exponential blow-up beginning around the 28-qubit mark.
Figure 4: Speedup as a function of the number of cores on a 24-core Intel Xeon w7-2495X processor for a 40-qubit LUCJ circuit containing 43 controlled-phase gates, with gate angles sampled from $\mathcal{N}(0,10^{-3})$ and total extent 1.732. For each data point, Raw Estimate was used to compute probabilities for 1,000 bitstrings and times were averaged over 10 trials.
Figure 5: Average runtime required to estimate a bitstring probability using Raw Estimate for a 100-qubit LUCJ circuit containing 111 controlled-phase gates, with gate angles drawn from $\mathcal{N}(0,10^{-3})$ and total circuit extent 5.281. Data points were collected for Hamming weights (number of electrons) ranging from 0 to 100 in increments of 2 and averaged over 10 trials. Simulations were performed on a 24-core Intel Xeon w7-2495X processor.
V-B Latency Benchmarks
In this section, we focus on the Raw Estimate algorithm, the fundamental subroutine of the Estimate and an approximate version of the Exact algorithm. Raw Estimate can be run with an arbitrary number of trajectories, making it more general and practically useful. At its core, a Raw Estimate computation amounts to summing contributions from a collection of sampled trajectories. Each trajectory corresponds to a particular decomposition of the controlled-phase gates and contributes an expectation value that must be accumulated into the overall estimate. Accordingly, we examine its runtime behavior, with attention not only to the number of trajectories required but also to non-asymptotic considerations that significantly impact the cost of evaluating each trajectory.
For a given trajectory $x$, the two main costs are constructing the mode transformation matrix $V(x)$ and evaluating the expectation value $\left\langle b\right|V(x)\left|a\right\rangle$. The latter is dominated by determinant calculations, which require selecting an $h\times h$ submatrix of $V(x)$, where $h$ is the Hamming weight of $\left\langle b\right|$ and $\left|a\right\rangle$. In the computational basis, $h$ corresponds to the number of set bits, or equivalently, the number of $X$ gates applied to prepare the Hartree-Fock state in the context of LUCJ circuits. Thus, states with larger Hamming weights—i.e., circuits with more fermions—incur higher computational cost. This can be seen clearly in Fig. 5, in which we fix the angles of a 100-qubit LUCJ circuit and compare the total time required to estimate the probability of a bitstring as a function of the Hamming weight. Independent of the number of trajectories, there is a clear trend indicating higher execution times for circuits for bitstrings with larger Hamming weights. In addition, Fig. 5 shows that for any given Hamming weight, when the number of trajectories is increased by a factor of 10, the respective runtime is also increased by approximately a factor of 10, as we would expect.
However, regardless of the Hamming weight of our bitstring, increasing the number of cores available to the simulator yields near-linear speedups. This is demonstrated in Fig. 4, which shows execution times for running Raw Estimate to estimate the probabilities of 1,000 bitstrings in parallel for different number of trajectories. Furthermore, we can see that as the task becomes harder (more trajectories are used), ExtraFerm scales more effectively.
V-C Error Analysis
Figure 6: The error between the true probability and probability obtained from Raw Estimate for all 400 outcome measurements in the support of a 12-qubit LUCJ circuit with 12 controlled-phase gates. These circuit correspond to fermionic systems with 6 spatial orbitals, 3 alpha spin orbitals, and 3 beta spin orbitals. They have extents of 3.753, 228.091, 2,254.514, 3,862.055, respectively (left to right). Orbital rotations are randomly initialized while controlled phase gate angles are sampled i.i.d. from $\sim\mathcal{N}(\theta,0.1)$. Results are displayed for $\theta\in\{0,\pi/3,2\pi/3,\pi\}$ across trajectory counts of 1,000, 10,000, 100,000, and 1,000,000. Each point is averaged over 10 trials. The top row displays absolute error $|\hat{p}-p_{\text{true}}|$ as a function of $p_{\text{true}}$ while the bottom row displays relative error $|\hat{p}-p_{\text{true}}|/p_{\text{true}}$ as a function of $p_{\text{true}}$.
Figure 7: Relative error of Raw Estimate versus trajectory count for a 40-qubit LUCJ circuit with 43 controlled-phase gates. These circuits correspond to fermionic systems with 20 spatial orbitals, 2 alpha spin orbitals, and 2 beta spin orbitals. They have extents of 3.484, 62.003, 1,216,201.946, and 8,766,956,281,978.891, respectively (left to right). Orbital rotations are randomly initialized while controlled phase gate angles are sampled i.i.d. from $\mathcal{N}(\theta,10^{-3})$ with $\theta\in\{\pi/64,\pi/16,\pi/4,\pi\}$. Each plot uses the median-probability bitstring from its respective circuit; the bitstring and true probability $p_{\text{true}}$ differ across panels. Blue markers show mean relative error over 10 trials. The horizontal black dashed line marks relative error $\epsilon_{\text{rel}}=.1$; the vertical magenta dashed line is the upper bound on the number of trajectories needed to achieve absolute error $\epsilon_{\text{abs}}=p_{\text{true}}\cdot.1$ with $\delta=10^{-4}$ (using $p_{\max}=p_{\text{true}}$). The green shading highlights the region where both conditions hold: $|\hat{p}-p_{\text{true}}|\leq p_{\text{true}}\cdot.1$ and the trajectory count is at or below the upper bound.
We now share error trends observed when using Raw Estimate to estimate probabilities for LUCJ circuits. We begin with a case study of a 12-qubit LUCJ circuit for which we estimated the probabilities of all outcome measurements in the support with different trajectory counts while varying the average controlled-phase gate angles. In Fig. 6, we display both the observed absolute and relative error for the entire support of the circuit. Across all circuits, we observe that absolute error is proportional to the magnitude of Born-rule probabilities while relative error is inversely proportional to the magnitude of Born-rule probabilities. The fact that a bitstring’s absolute error is proportional to its Born-rule probability ends up being very useful in practice. In section VI, we use this to successfully distinguish high-probability bitstrings from low-probability bitstrings with very small computational overhead.
For all four circuits, we observe decreases in both absolute and relative error as the number of trajectory counts increases. This trend is consistent with the theoretical guarantees of Raw Estimate, which guarantees decreasing additive error as the trajectory count grows. Comparisons between the four circuits are more nuanced. Varying the controlled-phase gate angles alters the Born-rule probabilities of the measurement outcomes. As these angles increase, the output distribution flattens: probability mass spreads across a larger fraction of the support, reducing the dominance of a few high-probability outcomes. This redistribution generally increases both absolute and relative errors. Thus, while increasing trajectory counts consistently reduces error across all circuits, larger controlled-phase angles shift the overall error profile upward.
This phenomenon is illustrated in Fig. 7, which reports the relative error as a function of trajectory count. We examine four 40-qubit LUCJ circuits that again differ only in the angles of their controlled-phase gates. To isolate the effect of these angles on error behavior, we select the median-probability bitstring from the support of each circuit and track its estimated error across increasing trajectory counts. The specific bitstring differs between circuits, but in each case we mark on the $x$-axis the theoretical trajectory count required to achieve a relative error of $0.1$, with the same threshold also indicated by a dashed line on the $y$-axis. For small gate angles ($\theta\approx 0$), this target is reached with relatively few trajectories. As $\theta$ increases, however, the observed errors remain greater than the desired level, and the error curves become steeper, demonstrating a larger difference in performance based on the number of trajectories used.
VI Applications
Figure 8: A schematic of our modification to the SQD algoritm. For the warm-start approach, the green path is taken after the first iteration of configuration recovery, and ExtraFerm is used to calculate Born-rule probabilities of bitstrings that have been corrected to the proper particle number. The gray path is taken for the subsequent iterations of the algorithm. Our experiments demonstrate that this modification improves the accuracy of ground-state energy calculations and reduces the variance between trials.
VI-A Sample-Based Quantum Diagonalization
Sample-Based Quantum Diagonalization (SQD) is an algorithm introduced in Ref. [robledo2025sqd] as an extension of Quantum-Selected Configuration Interaction [kanno2023quantum]. The central idea is to sample chemistry circuits on noisy quantum hardware and then apply classical post-processing to estimate molecular ground-state energies. A key property of LUCJ circuits is that, beyond conserving the total Hamming weight of a bitstring, they also preserve the Hamming weights of the $\alpha$- and $\beta$-spin sectors independently. Equivalently, an LUCJ state can be expressed as $\left|\Psi\right\rangle=\left|\beta\,\alpha\right\rangle$, where the spin sectors must retain their individual Hamming weights after the circuit is applied. SQD corrects bitstrings measured on quantum hardware that violate these conservation rules prior to using them in the energy calculation. Note that SQD is a variational algorithm, which guarantees that the energy estimates produced by it can never be lower than the ground-state energy. In other words, lower energy estimates are better.
We give a brief overview of SQD. First, a molecule is mapped to an LUCJ circuit $\left|\Psi\right\rangle$. Second, this circuit is sampled on a quantum computer to collect $\widetilde{\mathcal{X}}$, a set of noisy bitstrings representing electronic configurations. Third, bitstrings with incorrect particle number are then restored through an iterative procedure called configuration recovery to produce the set of bitstrings $\mathcal{X}_{\text{R}}=\mathcal{X}_{N}\cup\mathcal{X}_{\rightarrow N}$ where $\mathcal{X}_{N}$ and $\mathcal{X}_{\rightarrow N}$ are the bitstrings that originally had the correct particle number and the bitstrings that have been restored to have the correct particle number, respectively. Note that $\mathcal{X}_{\text{R}}$ is not a set in the mathematical sense; duplicate bitstrings are allowed. Fourth, $\mathcal{S}$ is subsampled from $\mathcal{X}_{\text{R}}$ according to observed bitstring frequencies, and the chemistry Hamiltonian is diagonalized in this subspace to obtain an energy value. Lastly, the average orbital occupancies are calculated, which are used to inform the diagonalization in the next iteration of SQD.
VI-B Warm-Start Approach
We propose a simple, cost-effective modification to the SQD workflow that we call warm-starting (see Fig. 8). For the original SQD workflow, $\mathcal{S}$ is subsampled from $\mathcal{X}_{\text{R}}$ according to the empirical frequencies of the bitstrings in $\mathcal{X}_{\text{R}}$. However, due to the large support of these circuits, each corrected bitstring in $\mathcal{X}_{\text{R}}$ appears extremely infrequently for nontrivial chemical systems. This means that $\mathcal{S}$ is subsampled approximately uniformly from $\mathcal{X}_{\text{R}}$, that is, each bitstring is chosen with a probability of $\approx\frac{1}{\#\text{ shots}}$. Since SQD is a variational algorithm, this method of subsampling is beneficial, allowing the solver to explore a diverse set of electronic configurations and converge to the ground state energy.
However, average orbital occupancies identified during early iterations of SQD inform the Hamiltonian diagonalization for later iterations of SQD. Since these bitstrings are chosen randomly, energy estimates can suffer from high variance between trials. Our warm-start approach addresses this by using ExtraFerm to calculate the probabilities of the bitstrings in $\mathcal{X}_{\text{R}}$ during the first iteration of configuration recovery and then choosing the bitstrings with the highest probabilities, rather than subsampling. Later iterations of SQD proceed as usual, subsampling from $\mathcal{X}_{\text{R}}$ according to the empirical frequencies of the corrected bitstrings. We find that the warm-start approach produces energy estimates that are on average more accurate than the ones produced by the original SQD workflow and have lower variance across trials.
VI-C Experimental Setup
[…]
For both systems, probability estimation with ExtraFerm used to warm-start SQD was performed using the Raw Estimate algorithm with 1,000 trajectories per bitstring. In the following section, we report the ExtraFerm overhead as the added percentage to the baseline SQD runtime incurred running Raw Estimate for all sampled bitstrings. This cost is independent of the subspace dimension. Finally, all computations in this section were performed on 12 cores of an Intel Gold 6248R processor.
[…]
Subspace Dim Avg. Error Reduction Avg. Var. Red. Med. Var. Red. Runtime ExtraFerm Overhead
$200^{2}=40,000$ $60.15\%$ $71.06\%$ $90.32\%$ 37m $4.07\%$
$300^{2}=90,000$ $59.74\%$ $92.10\%$ $98.29\%$ 38m $4.00\%$
$400^{2}=160,000$ $52.29\%$ $92.79\%$ $97.54\%$ 40m $3.71\%$
$500^{2}=250,000$ $43.56\%$ $79.05\%$ $87.59\%$ 44m $3.41\%$
[…]
Subspace Dim Avg. Error Reduction Avg. Var. Red. Med. Var. Red. Runtime ExtraFerm Overhead
$500^{2}=250,000$ $37.63\%$ $82.24\%$ $87.76\%$ 1h 22m $2.03\%$
$1,000^{2}=1,000,000$ $46.09\%$ $33.48\%$ $98.34\%$ 2h 47m $1.00\%$
$1,500^{2}=2,250,000$ $32.78\%$ $44.49\%$ $65.88\%$ 4h 28m $0.62\%$
$2000^{2}=4,000,000$ $36.24\%$ $44.66\%$ $73.61\%$ 6h 40m $0.42\%$
[…]
VI-D Classical Simulation of a Hydrogen Chain
[…]
These LUCJ circuits were small enough to be sampled classically using ffsim [ffsim]. We sampled 100,000 bitstrings from each circuit with depolarizing noise set to $0.9$ to model realistic quantum hardware. For this system, we studied subspace dimensions of $\{200^{2},300^{2},400^{2},500^{2}\}$. On average, across all four plots, we observe reductions in both error and variance for warm-started SQD. As shown in Fig 9 and Table I, the reductions in error ranged between $43.56\%$ and $60.15\%$ relative to the full configuration interaction (FCI) energy. The average variance reductions ranged between $71.06\%$ and $92.79\%$. We also report median variance reductions, which range between $87.59\%$ and $98.29\%$. The additional cost of employing ExtraFerm to warm-start SQD was minimal. On average, it required 90 seconds per energy curve, corresponding to an added overhead of at worst $4.07\%$ and at best $3.41\%$ with respect to the original SQD runtime.
VI-E Nitrogen on the Heron Quantum Processor
[…]
These LUCJ circuits were sampled on IBM’s Heron superconducting quantum processor. For this system, we tested subspace dimensions of $\{500^{2},1000^{2},1500^{2},2000^{2}\}$. On average, across all four plots, we again observe reductions in both error and variance for warm-started SQD. As shown in Fig. 9 and Table I, the reductions in error ranged between $32.78\%$ and $46.09\%$ relative to the heat-bath configuration interaction (HCI) energy. The average variance reductions ranged between $33.48\%$ and $82.24\%$, and median variance reductions range between $65.88\%$ and $98.34\%$. In this case, the additional cost of employing ExtraFerm to warm-start SQD was negligible. On average, it required 100 seconds per energy curve, corresponding to an added overhead of at worst $2.03\%$ and at best $42\%.$
VI-F Summary
We observe that ExtraFerm offers substantial performance benefits to SQD through the warm-start approach with very little computational cost. For both molecules, across all subspace dimensions, we see that on average, the warm-started variant of SQD obtained more accurate ground-state energy estimates with lower variance between trials. Most importantly, these results highlight a practical strategy for enhancing the utility of near-term quantum devices for chemistry applications.
[…]
VII Related Works
VII-A Clifford+T Circuits
The Clifford+T gate set, comprising the Clifford group (generated by the phase, Hadamard, and CNOT gates) and the T gate, is among the most extensively studied in quantum computing. Stabilizer circuits, which contain only Clifford gates, are classically simulable in polynomial time by the Gottesman-Knill theorem [gottesman1998heisenberg]. Adding T gates enables universal quantum computation [bravyi2005universal]. States produced by a small number of T gates have low stabilizer rank and can be expressed as a concise linear combination of stabilizer states [bravyi2019simulation], making them amenable to classical simulation. Ref. [bravyi2016improved] leverages this property to propose a Monte Carlo probability estimation routine for Clifford-dominated circuits. This algorithm is similar in spirit to Reardon-Smith’s [reardon2024improved], which ExtraFerm builds upon.
VII-B Tensor Networks
Tensor networks [white1992density, orus2014practical] are a framework for representing quantum many-body systems and are a popular choice for simulating quantum circuits [markov2008simulating]. Tensor networks factorize quantum states into smaller tensors, which works particularly well in instances for which there is limited entanglement. However, this method may scale poorly for some quantum circuits with many nonlocal gates.
VII-C Classical Simulation to Boost Quantum Fidelity
Recently, several works have used classical simulation to mitigate errors on quantum computers. These methods typically simulate a subset of the quantum system or an efficiently simulable related system. For instance, Clifford Data Regression [czarnik2021error] fits a linear model using Clifford-dominated circuits to map noisy observables to their noiseless counterparts, achieving error reductions between one and two orders of magnitude. Liu et al. [liu2022classical] offload the computation of an auxiliary error-mitigation circuit to a classical processor, further reducing noise effects. Finally, CAFQA [ravi2022cafqa] accelerates convergence of variational quantum algorithms by searching over Clifford circuits to identify high-quality initial parameters, yielding substantial accuracy gains.
VIII Discussion
As the utility of near-term quantum computers increases, researchers need tools to assess the quality of data collected on noisy quantum hardware. We have introduced ExtraFerm, a simulator that can be used to estimate Born-rule probabilities of bitstrings with respect to particle number-conserving extended matchgate circuits. We have seen that, in addition to being a useful tool for simulating chemistry circuits, ExtraFerm can be integrated with SQD to improve the quality of molecular ground-state energy estimates through a warm-start technique. We propose two directions for future work in which ExtraFerm could further advance the practical application of quantum chemistry on current quantum hardware.
First, ExtraFerm could enable sophisticated post-processing strategies for noisy samples from quantum hardware. While we found that the simple approach of selecting high-probability bitstrings during early iterations of SQD improved energy estimates, there are a variety of other techniques that one could try. The ability to compute Born-rule probabilities of existing data could provide a way to re-weight the observed bitstrings and re-sample from the pool of collected data. Discrepancies between the empirical and simulated distributions could also be used to infer or correct aspects of the underlying noise model of the device. In this way, ExtraFerm may support powerful error-mitigation strategies for chemistry circuits on near-term quantum devices.
Second, rather than using ExtraFerm to post-process measurements obtained from noisy quantum devices, one could employ it to optimize the parameters of a quantum circuit before it is executed on hardware. For instance, in a quantum chemistry setting, suppose one has an approximate ansatz that prepares a target electronic state. Given access to estimates of amplitudes in the target state, ExtraFerm could be used to minimize the discrepancy between the target probabilities of selected bitstrings (corresponding to important configurations) and the probabilities produced by the circuit. Gradient-free methods such as simultaneous perturbation stochastic approximation (SPSA) [spall2002implementation] could be used to facilitate optimization.
We hope that researchers working with matchgate-based chemistry circuits will find ExtraFerm useful and discover new ways of incorporating it into their workflows. Furthermore, we hope that the spirit of our integration of ExtraFerm with SQD will encourage the development of more hybrid approaches that use classical computation to extend the capabilities of quantum chemistry on near-term quantum hardware.
Acknowledgments
The calculations in Section VI were performed on the Midway3 Cluster, maintained by the University of Chicago Research Computing Center. We obtained the nitrogen dataset of Ref. [robledo2025sqd] from https://github.com/jrm874/sqd_data_repository.
-A Proof of Exact Calculation
$p_{\text{true}}$ $=\left|\left\langle b\right|V_{k+1}\prod_{j=1}^{k}C(\theta_{j})V_{j}\left|a\right\rangle\right|^{2}$ (20)
$=\left|\left\langle b\right|V_{k+1}\prod_{j=1}^{k}\left(\cos\left(\frac{\theta_{j}}{4}\right)D_{0}(\theta_{j})+i\sin\left(\frac{\theta_{j}}{4}\right)D_{1}(\theta_{j})\right)V_{j}\left|a\right\rangle\right|^{2}$ (21)
$=\left|\left\langle b\right|\sum_{x\in\{0,1\}^{k}}V_{k+1}\prod_{j=1}^{k}\left(\cos\left(\frac{\theta_{j}}{4}\right)D_{0}(\theta_{j})\right)^{x_{j}-1}\cdot\left(i\sin\left(\frac{\theta_{j}}{4}\right)D_{1}(\theta_{j})\right)^{x_{j}}V_{j}\left|a\right\rangle\right|^{2}$ (22)
$=\left|\sum_{x\in\{0,1\}^{k}}\left(\prod_{j=1}^{k}\cos\left(\frac{\theta_{j}}{4}\right)^{x_{j}-1}\cdot\left(i\sin\left(\frac{\theta_{j}}{4}\right)\right)^{x_{j}}\right)\left\langle b\right|V(x)\left|a\right\rangle\right|^{2}$ (23)